OpenStudy (anonymous):

Help 1 question Step by step

OpenStudy (anonymous):

Help anyone who can help me step by step

OpenStudy (jdoe0001):

\(\bf 2x^2=-3x-5\implies 2x^2+3x+5=0\\ \quad \\ \textit{use the quadratic formula to get the values for "x"}\\ \quad \\ \text{quadratic formula}\\ x= \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\)

OpenStudy (anonymous):

yea can you solve it for me steo by step @jdoe0001

OpenStudy (jdoe0001):

\( \begin{array}{llll} a&b&c\\ 2x^2&+3x&+5=0 \end{array}\\ \quad \\ \bf \text{quadratic formula}\\ x= \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\implies x= \cfrac{ - (3) \pm \sqrt { (3)^2 -4(2)(5)}}{2(2)}\\ \quad \\ x= \cfrac{ - 3 \pm \sqrt { 9 -40}}{4}\implies x= \cfrac{ - 3 \pm \sqrt { -31}}{4} \)

OpenStudy (anonymous):

okay thanks :)

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

i have 5 more questions can you help me with the others?

OpenStudy (anonymous):

i have answers just need to know if they are right?

OpenStudy (jdoe0001):

I'll be dashing in a couple of mins

OpenStudy (anonymous):

okay here they are 1.b 2.a 3.d 4.c 5.c

OpenStudy (anonymous):

@jdoe0001 here they are above?

OpenStudy (jdoe0001):

pretty straightforward, just use the quadratic formula on all, once you set them in the standard form, that is = 0, just like above

OpenStudy (anonymous):

okay thanks